1.3 cm Session
Session: Feb 2004 (session 1)
Wavelength: 1.3 cm
To give feedback on this session, click
1.3 cm Session Feedback
Medicina (Z. Paragi at Mon Feb 16 12:56:48 2004) Failure :
The LO was not locked to the right frequency.
Urumqi (Z. Paragi at Mon Feb 16 13:51:21 2004) Failure :
The LO was locked to 22100 MHz instead of 22000 MHz before N04K1.